In a trial of a cytotoxic drug for lung cancer treatment, the investigators reported 6 deaths among 60 patients treated with this agent, compared with 15 deaths among 58 controls on placebo. Calculate the attributable benefit of this agent in reducing lung cancer-related mortality in this population.
Answer #1

Cumulative incidence in the unexposed group (control group), CIu = 15 / 58 = 0.2586207

Cumulative incidence in the exposed group (agent group), CIe = 6 / 60 = 0.1

Attributable benefit of this agent = (CIu - CIe) / CIu

= (0.2586207 - 0.1) / 0.2586207

= 0.6133333

= 61.33%